The reaction, A → products, is known to be zeroth order with respect to the concentration of A and zeroth order overall. The rate constant for the reaction is 0.0750 M·hr-1 at a certain temperature.

If the initial concentration of A is 0.875 M. What will be the concentration of A after 10.0 hr?
Answer M

How long would it take for 80.0% of the starting material (A) to undergo reaction?
Answer hr
Answer #1

Since the reaction is following zero order kinetics .

So, we have the expression :

k = { [A]• - [A] } / t

Where t is time and [A]• is initial concentration

[A] is final or remaining concentration of A

And k is rate constant .

So, k = 0.0750 Mhr^-1 and [A]• = 0.875 M

t = 10 hr

So, [A] = 0.875 - ( 0.0750 X 10 ) = 0.125 M i.e remaining concentration after 10 hr

and from statement that 80% of starting material undergo reaction means of we have 100% initially then remaining amount is 20% so,

t = 100 - 20 / 0.0750 = 1066.67 hr